# shapes,patterns,positioning

### v2

@@ Abstract
[2]organization=Gothenburg University, country=Sweden
@@ 1 Introduction
In this paper we make several contributions to the development of simplicial homotopy theory, in particular the Kan-Quillen model structure, in an explicit or constructive style. Our motivation comes from homotopy type theory Uni (13), where Kan fibrations have been used in KL (21) to model dependent types for homotopy type theory. An essential part of this proof is that Kan fibrations are closed under pushforwards, so we can model dependent products.
Homotopy type theory in turn models constructive mathematics. However, in BCP (15) (see also Par (18)), it has been shown that a proof that Kan fibrations are closed under pushforwards cannot be made constructively. This problem has been addressed by moving away from simplicial sets, and instead use cubical sets, as in CCHM (18) and BCH (14). However, as simplicial sets are important and pervasive in homotopy theory, it is still desirable to have a constructive notion of Kan fibrations in simplicial sets.
At this point in time, there are two notions of Kan fibrations in simplicial sets which can constructively be shown to be closed under pushforwards, namely the effective Kan fibrations from BF (22) and the uniform Kan fibrations from GS (17). We will not discuss uniform Kan fibration here, but let us note that this notion of fibration is not “local”, as discussed in Appendix D of BF (22), and every effective Kan fibration is also a uniform Kan fibration, as shown in Corollary 11.2 of BF (22).
Therefore we will focus on the effective Kan fibrations, which are Kan fibration equipped with explicit solutions to lifting problems against horns satisfying a certain compatibility condition. The hope is that for these effective Kan fibrations we can obtain constructive and explicit versions of many standard results in simplicial homotopy theory, as was done in BF (22) for closure under pushforwards. In this paper we will show that that is indeed the case for two other standard results.
The first results says that simplicial groups are Kan complexes Moo (58). The proof of this result is constructive, in that the argument constructs explicit lifts again horn inclusion. We will show that these lifts satisfy the additional compatibility requirement making them effective Kan complexes. In fact, we will show that all simplicial Malcev algebras are effective Kan complexes, where Malcev algebras are exactly those algebras for which all simplicial objects are Kan complexes, as shown in CKP (93) and JP (02). We will describe these results in Section 3.
The second result is that Kan fibrations are the right class in a weak factorisation system. By now there is a well-established explicit version of a weak factorisation system, which goes by the name of an “algebraic weak factorisation system” Gar (08); BG (16); Bou (23). In Section 4 we will show that the effective Kan fibrations are the right class in a algebraic weak factorisation system.
@@ 2.1 Mathematical background on Kan fibrations
In this subsection, we shall recall some concepts from higher category theory. We assume some knowledge on category theory and topos theory. References include Lur (23), Rie (16) and Joh (77).
@@
    For $0\leq i\leq n$, we have a face map ${d_{i}}$ skipping over $i$.
@@
Every map in $\Delta$ can be written as composition of face and degeneracy maps. Between the face and degeneracy maps, we have the following composition laws, called simplicial identities:
@@
A simplicial set is a presheaf on $\Delta$. More generally, for any category $\mathcal{C}$, a simplicial object of $\mathcal{C}$ is a functor $\Delta^{op}\to\mathcal{C}$. The category of simplicial sets is denoted sSet. The representable presheaf corresponding to $[n]$ is denoted ${\Delta^{n}}$ and called the $n$-simplex. For a simplicial set $G$, the Yoneda lemma allows us to identify maps $\Delta^{n}\to G$ as elements of $G_{n}$.
@@
A simplicial set $X$ is called a Kan complex if the unique map to the terminal object $X\to\Delta^{0}$ is a Kan fibration. We will use the same vocabulary for the other notions of Kan fibrations introduced in this paper.
@@ 2.2 Degenerate-preferring Kan fibrations
A functional Kan fibration is a morphism of simplicial sets $\alpha:X\to Y$, together with a function lift, which takes as input maps $x,y$ making the solid diagram in Equation 7 commute, and outputs a morphism $\text{lift}(x,y):\Delta^{n}\to X$ making the diagram commute.
The definition of the degenerate-preferring Kan fibrations can be motivated by the following result of Van den Berg, Faber and Sattler. In Appendix C of BF (22) they show that if the lifting problem in Equation 7 has a degenerate solution, it is unique:
@@
Suppose that both $z\circ s_{j}$ and $z^{\prime}\circ s_{j^{\prime}}$ are possible lifts for lifting problems as in Equation 7. Then $z\circ s_{j}=z^{\prime}\circ s_{j^{\prime}}$.
@@
Evan Cavello has pointed out to us that degenerate-preferring Kan fibrations are similar to the “regular” or “normal” Kan fibrations in cubical sets (see for example the introduction of Swa (18)). Thierry Coquand based this name on the regular fiber spaces, which first appeared in Hur (55).
@@ 2.3 Symmetric effective Kan fibrations
Effective Kan fibrations are functional Kan fibrations as in Definition 2.1 where the function lift satisfies certain conditions. Inspired by Theorem 12.1 from BF (22), these conditions amount to stability along pullback along degeneracy maps. We will first give an example of what this means for a specific lifting problem, and then give the general definition.
@@
For the general case, consider Equation 7, and suppose we pull back $\iota$ along $s_{j}:\Delta^{n+1}\to\Delta^{n}$:
[…]
We shall argue that the inclusion $s_{j}^{*}(\Lambda^{n}_{m})\hookrightarrow\Delta^{n+1}$ factors through a horn inclusion $\iota^{*}:\Lambda^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$. We shall then show that a dashed arrow as above allows us to extend $x\circ\sigma$ to a map $s_{j}^{*}(x):\Lambda^{n+1}_{m^{*}}\to G$.
@@
If there exists a map $\text{lift}(x,y):\Delta^{n}\to X$ making the right-most square above commute, we could extend $x\circ\sigma$ to a map $s_{j}^{*}(\Delta^{n}_{m})\cup d_{j_{\pm}}\to X$. which has $\text{lift}(x,y)$ as value on the face $d_{j_{\pm}}$. As the intersection of $d_{j_{\pm}}$ and $s_{j}^{*}(\Lambda^{n}_{m})$ is given by $\Lambda^{n}_{m}$, where $\text{lift}(x,y)$ equals $x$, this is properly defined. This allows us to extend $x\circ\sigma$ to all of $\Lambda^{n+1}_{m^{*}}$.
@@
So to summarize, if we have a dashed arrow $\text{lift}(x,y)$ in Equation 7, for all $m^{*}$ with
@@ 2.4 Effective Kan fibrations
In this section we recall the definition of effective Kan fibrations. The definition we give here is based on Theorem 12.1 in BF (22). We give a slightly different presentation here which is more in line with the rest of this paper.
@@
The argument in Corollary 12.2 of BF (22) shows that degenerate-preferring Kan fibrations must be effective Kan fibrations. In Section 3, we shall show constructively that a large class of Kan fibrations can constructively be given the structure of degenerate-preferring Kan fibrations. Therefore, they are also effective Kan fibrations.
@@ 3 Simplicial Malcev algebras
A well-known result of Moore (Moo, 58, Theorem 3.4) says that simplicial groups are Kan complexes. Since then, this result has been generalized. In CKP (93) it was shown that we can classify all algebras for which the simplicial objects are Kan complexes, namely the Malcev algebras. In addition, in JP (02) it is shown that any surjection of simplicial Malcev algebras is a Kan fibration.
In this section, we shall generalize the latter result to degenerate-preferring Kan fibrations. First we shall recall some context on simplicial Malcev algebras and introduce some new terminology.
@@ 3.1 Malcev algebras
In JP (02), Proposition 3, a concise and constructive proof is given that implies the following:
@@
So if we have a constructive proof that all simplicial algebras of $\mathcal{T}$ are Kan we have a constructive proof that $\mathcal{T}$ is Malcev. The converse of this theorem is also known to be true, and was even generalized further in JP (02), Theorem 4 to the following:
@@ 3.2 A degenerate-preferring structure
To show the above Theorem, we will first recall in Construction 3.6 the lifting method that JP (02) uses to solve the required lifting problem. We will then show two properties of this lifting structure and conclude it is degenerate-preferring.
@@
Let $\alpha:X\to Y$ be a morphism of simplicial Malcev algebras with degeneracy-section $(\beta_{n}:Y_{n}\to X_{n})_{n\in\mathbb{N}}$. Now consider maps $x,y$ as in Equation 7.
@@
In Theorem 4 of JP (02) the following is shown:
@@
Suppose that $X,Y$ are Malcev algebras and $\alpha$ has a degeneracy-section in Equation 7. Then the map $\mathrm{lift}(x,y)$ from Construction 3.6 solves the lifting problem.
@@
For the rest of this section, we will say that $k$ is encountered before $l$ or $l$ is encountered after $k$ is $w_{k}$ is defined before $w_{l}$ in construction 3.6. With this terminology, we can formulate the rest of our proof strategy
@@
For a lifting problem with a degenerate solution as in Equation LABEL:eqnLiftProbWithDegSol and $w_{k}$ as in Construction 3.6, whenever $w_{k}$ is defined before $w_{j}$ and $w_{j+1}$, we have $w_{k}=z\circ s_{j}$ for some $z\in X_{n}$.
@@
Note that $x\circ d_{k}=g\circ s_{j}\circ d_{k}$ by Equation LABEL:eqnLiftProbWithDegSol. We will now calculate $s_{j}\circ d_{k}\circ s_{k^{\prime}}$. As $k$ is encountered before $j,j+1$, we have $k\neq j,j+1$. We will now make a case distinction on whether $k<j$ or $k>j+1$.
@@
For any $k$, a solution to the lifting problem in Equation 7 is a fixed point for $N_{k}$.
@@
So to summarize, to show Theorem 3.5, we use Construction 3.6 to give a lifting structure for $\alpha$. It is really a lifting structure by Theorem 4 for JP (02). We showed that if a degenerate solution of the form $g\circ s_{j}$ exists, the helper functions $w_{k}$ from Construction 3.6 will always be degenerate along $j$, and as soon as $k$ is encountered within or after $\{j,j+1\}$, we have $w_{k}=g\circ s_{j}$. In particular $\text{lift}(x,y\circ s_{j})=g\circ s_{j}$. Therefore the lifting structure for $\alpha$ is degenerate-preferring.
@@ 3.3 Some consequences
Let $X$ be a inhabited simplicial algebra of a Malcev theory. $X$ is inhabited iff $X_{0}$ is inhabited. Let $x\in X_{0}$ witness this. Then the morphism $\beta:\Delta^{0}\to X$ given by $\beta_{n}(*)=(s_{0})^{n}(x)$ is a section, and hence a degeneracy-section of the unique morphism $X\to\Delta^{0}$.
@@
Corollary 3.12 says that we have found all algebraic theories $\mathcal{T}$ such that simplicial algebras of $\mathcal{T}$ are degenerate-preferring Kan. However, we could look further than simplicial algebras. There is also a notion of Malcev categories, which are regular categories with a condition on the lattice of relations. In Mal (54), Malcev222 This reference is in Russian, and the authors were unable to find a translation. The name Malcev is a transcription, and has also been transcribed as Mal’cev, Maltsev and Mal’tsev. showed that for algebras, this is equivalent to having a Malcev operation. Let us mention Theorem 4.6 of CKP (93) which states that a regular category $\mathcal{A}$ is a Malcev category iff every simplicial object of $\mathcal{A}$ is a Kan complex.
@@ 4 Algebraic weak factorisation systems
The classical Kan fibrations are part of the “right class” in the weak factorisation system of the Kan-Quillen model structure. In this section, we will discuss a constructive variation on weak factorisation systems, namely that of an algebraic weak factorisation system (awfs). We will show that the symmetric effective and the effective Kan fibrations both fit into their own awfs. In the first two subsections, we will recall the definition of a double category and an awfs; the definition of an awfs we give here is the one from Bou (23), where it is called a lifting algebraic weak factorisation system. Indeed, these sections are based on Bou (23) and contain no original work. We will then define a class of double categories based on horn inclusions and study awfss where the left double category comes from this class. We will then characterize the vertical morphisms of the right double category of such an awfs and show that the effective and symmetric effective Kan fibrations satisfy this characterization.

### v3

# Examples and cofibrant generation of effective Kan fibrations

@@ Abstract
[2]organization=Department of Computer Science and Engineering
University of Gothenburg and Chalmers University of Technology, country=Sweden
@@ 1 Introduction
In this paper we make several contributions to the development of simplicial homotopy theory, in particular the Kan-Quillen model structure, in an explicit or constructive style. Our motivation comes from homotopy type theory [23], where Kan fibrations have been used in [15] to model dependent types for homotopy type theory. An essential part of this proof is that Kan fibrations are closed under pushforwards, so we can model dependent products.
Homotopy type theory in turn models constructive mathematics. However, in [2] (see also [19]), it has been shown that a proof that Kan fibrations are closed under pushforwards cannot be made constructively. This problem has been addressed by moving away from simplicial sets, and instead use cubical sets, as in [6] and [1]. However, as simplicial sets are important and pervasive in homotopy theory, it is still desirable to have a constructive notion of Kan fibrations in simplicial sets.
At this point in time, there are two notions of Kan fibrations in simplicial sets which can constructively be shown to be closed under pushforwards, namely the effective Kan fibrations from [3] and the uniform Kan fibrations from [11]. We will not discuss uniform Kan fibration here, but let us note that this notion of fibration is not “local”, as discussed in Appendix D of [3], and every effective Kan fibration is also a uniform Kan fibration, as shown in Corollary 11.2 of [3].
Therefore we will focus on the effective Kan fibrations, which are Kan fibrations equipped with explicit solutions to lifting problems against horns satisfying a certain compatibility condition. The hope is that for these effective Kan fibrations we can obtain constructive and explicit versions of many standard results in simplicial homotopy theory, as was done in [3] for closure under pushforwards. In this paper we will show that that is indeed the case for two other standard results.
The first results says that simplicial groups are Kan complexes [18]. The proof of this result is constructive, in that the argument constructs explicit lifts against horn inclusions. We will show that these lifts satisfy the additional compatibility requirement making them effective Kan complexes. In fact, we will show that all simplicial Malcev algebras are effective Kan complexes, where Malcev algebras are exactly those algebras for which all simplicial objects are Kan complexes, as shown in [7] and [14]. We will describe these results in Section 3.
The second result is that Kan fibrations are the right class in a weak factorisation system. By now there is a well-established explicit version of a weak factorisation system, which goes by the name of an “algebraic weak factorisation system” [8, 4, 5]. In Section 4 we will show that the effective Kan fibrations are the right class in an algebraic weak factorisation system.
@@ 2.1 Mathematical background on Kan fibrations
In this subsection, we shall recall some concepts from higher category theory. We assume some familiarity with category theory and topos theory. References include [16], [20] and [13].
@@
    For any $n\in\mathbb{N}$, $0\leq i\leq n$, we have a face map ${d_{i}}$ skipping over $i$.
@@
Every map in $\Delta$ can be written as a composition of face and degeneracy maps. Between the face and degeneracy maps, we have the following composition laws, called the simplicial identities:
@@
A simplicial set is a presheaf on $\Delta$. More generally, for any category $\mathcal{C}$, a simplicial object of $\mathcal{C}$ is a functor $\Delta^{op}\to\mathcal{C}$. The category of simplicial sets is denoted sSet. The representable presheaf corresponding to $[n]$ is denoted ${\Delta^{n}}$ and called the $n$-simplex. For a simplicial set $G$, the Yoneda lemma allows us to identify maps $\Delta^{n}\to G$ with elements of $G_{n}$.
@@
A simplicial set $X$ is called a Kan complex if the unique map to the terminal object $X\to\Delta^{0}$ is a Kan fibration. We will use the same vocabulary for the other notions of Kan fibration introduced in this paper.
@@ 2.2 Degenerate-preferring Kan fibrations
A functional Kan fibration is a morphism of simplicial sets $\alpha:X\to Y$, together with a function lift, which takes as input maps $x,y$ making the solid diagram in Diagram 7 commute, and outputs a morphism $\text{lift}(x,y):\Delta^{n}\to X$ making the diagram commute.
The definition of the degenerate-preferring Kan fibrations can be motivated by the following result of Van den Berg, Faber and Sattler. In Appendix C of [3] they show that if the lifting problem in Diagram 7 has a degenerate solution, it is unique:
@@
Suppose that both $z\circ s_{j}$ and $z^{\prime}\circ s_{j^{\prime}}$ are possible lifts for lifting problems as in Diagram 7. Then $z\circ s_{j}=z^{\prime}\circ s_{j^{\prime}}$.
@@
Evan Cavallo has pointed out to us that degenerate-preferring Kan fibrations are similar to the “regular” or “normal” Kan fibrations in cubical sets (see, for example, the introduction of [21] or [10, Section 7]). Thierry Coquand based this name on the regular fiber spaces, which first appeared in [12].
@@ 2.3 Symmetric effective Kan fibrations
Effective Kan fibrations are functional Kan fibrations as in Definition 2.1 where the function lift satisfies certain conditions. Inspired by Theorem 12.1 from [3], these conditions amount to stability along pullback along degeneracy maps. We will first give an example of what this means for a specific lifting problem, and then give the general definition.
@@
00111122223333 (11)
For the general case, consider Diagram 7, and suppose we pull back $\iota$ along $s_{j}:\Delta^{n+1}\to\Delta^{n}$:
@@
We shall argue that the inclusion $s_{j}^{*}(\Lambda^{n}_{m})\hookrightarrow\Delta^{n+1}$ factors through a horn inclusion $\iota^{*}:\Lambda^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$. We shall then show that a dashed arrow as above allows us to extend $x\circ\sigma_{j}$ to a map $s_{j}^{*}(x):\Lambda^{n+1}_{m^{*}}\to X$.
@@
If there exists a map $\text{lift}(x,y):\Delta^{n}\to X$ making the right-most square above commute, we could extend $x\circ\sigma$ to a map $s_{j}^{*}(\Lambda^{n}_{m})\cup d_{j_{\pm}}\to X$. which has $\text{lift}(x,y)$ as value on the face $d_{j_{\pm}}$. As the intersection of $d_{j_{\pm}}$ and $s_{j}^{*}(\Lambda^{n}_{m})$ is given by $\Lambda^{n}_{m}$, where $\text{lift}(x,y)$ equals $x$, this is properly defined. This allows us to extend $x\circ\sigma$ to all of $\Lambda^{n+1}_{m^{*}}$.
@@
So to summarize, if we have a dashed arrow $\text{lift}(x,y)$ in Diagram 7, for all $m^{*}$ with
@@ 2.4 Effective Kan fibrations
In this section we recall the definition of an effective Kan fibrations. The definition we give here is based on Theorem 12.1 in [3]. We give a slightly different presentation here which is more in line with the rest of this paper.
@@
The argument in Corollary 12.2 of [3] shows that degenerate-preferring Kan fibrations must be effective Kan fibrations. In Section 3, we shall show constructively that a large class of Kan fibrations can constructively be given the structure of degenerate-preferring Kan fibrations. Therefore, they are also effective Kan fibrations.
@@ 3 Simplicial Malcev algebras
A well-known result of Moore [18, Theorem 3.4] says that simplicial groups are Kan complexes. Since then, this result has been generalized. In [7] it was shown that we can classify all algebras for which the simplicial objects are Kan complexes, namely the Malcev algebras. In addition, in [14] it is shown that any surjection of simplicial Malcev algebras is a Kan fibration.
In this section, we shall generalize the latter result to degenerate-preferring Kan fibrations. First we shall give some context for simplicial Malcev algebras and introduce some new terminology.
@@ 3.1 Malcev algebras
In [14, Proposition 1] a concise and constructive proof is given that implies the following:
@@
So if we have a constructive proof that all simplicial algebras of $\mathcal{T}$ are Kan, we have a constructive proof that $\mathcal{T}$ is Malcev. The converse of this theorem is also known to be true, and was even generalized further in [14, Theorem 3] to the following:
@@ 3.2 A degenerate-preferring structure
To show the above theorem, we will first recall in Construction 3.6 the lifting method that [14] uses to solve the required lifting problem. We will then show two properties of this lifting structure and conclude it is degenerate-preferring.
@@
Let $\alpha:X\to Y$ be a morphism of simplicial Malcev algebras with degeneracy-section $(\beta_{n}:Y_{n}\to X_{n})_{n\in\mathbb{N}}$. Now consider maps $x,y$ as in Diagram 7.
@@
In the proof of Theorem 3 of [14] the following is shown:
@@
Suppose that $X,Y$ are Malcev algebras and $\alpha$ has a degeneracy-section in Diagram 7. Then the map $\mathrm{lift}(x,y)$ from Construction 3.6 solves the lifting problem.
@@
For the rest of this section, we will say that $k$ is encountered before $l$ or $l$ is encountered after $k$ is $w_{k}$ is defined before $w_{l}$ in construction 3.6. With this terminology, we can formulate the rest of our proof strategy:
@@
For a lifting problem with a degenerate solution as in Equation 22 and $w_{k}$ as in Construction 3.6, whenever $w_{k}$ is defined before $w_{j}$ and $w_{j+1}$, we have $w_{k}=z\circ s_{j}$ for some $z\in X_{n}$.
@@
Note that $x\circ d_{k}=g\circ s_{j}\circ d_{k}$ by Equation 22. We will now calculate $s_{j}\circ d_{k}\circ s_{k^{\prime}}$. As $k$ is encountered before $j,j+1$, we have $k\neq j,j+1$. We will now make a case distinction on whether $k<j$ or $k>j+1$.
@@
For any $k$, a solution to the lifting problem in Diagram 7 is a fixed point for $N_{k}$.
@@
So to summarize, to show Theorem 3.5, we use Construction 3.6 to give a lifting structure for $\alpha$. It is really a lifting structure by Theorem 3.7. We showed that if a degenerate solution of the form $g\circ s_{j}$ exists, the helper functions $w_{k}$ from Construction 3.6 will always be degenerate along $j$, and as soon as $k$ is encountered within or after $\{j,j+1\}$, we have $w_{k}=g\circ s_{j}$. In particular $\text{lift}(x,y\circ s_{j})=g\circ s_{j}$. Therefore the lifting structure for $\alpha$ is degenerate-preferring.
@@ 3.3 Some consequences
Let $X$ be a inhabited simplicial algebra of a Malcev theory. Then $X$ is inhabited iff $X_{0}$ is inhabited. Let $x\in X_{0}$ witness this. Then the morphism $\beta:\Delta^{0}\to X$ given by $\beta_{n}(*)=(s_{0})^{n}(x)$ is a section, and hence a degeneracy-section of the unique morphism $X\to\Delta^{0}$.
@@
Corollary 3.12 says that we have found all algebraic theories $\mathcal{T}$ such that simplicial algebras of $\mathcal{T}$ are degenerate-preferring Kan. However, we could look further than simplicial algebras. There is also a notion of Malcev categories, which are regular categories with a condition on the lattice of relations. In [17], Malcev222 This reference is in Russian, and the authors were unable to find a translation. The name Malcev is a transcription, and has also been transcribed as Mal’cev, Maltsev and Mal’tsev. showed that for algebras, this is equivalent to having a Malcev operation. Let us mention Theorem 4.6 of [7] which states that a regular category $\mathcal{A}$ is a Malcev category iff every simplicial object of $\mathcal{A}$ is a Kan complex.
@@ 4 Algebraic weak factorisation systems
The classical Kan fibrations are part of the “right class” in the weak factorisation system of the Kan-Quillen model structure. In this section, we will discuss a constructive variation on weak factorisation systems, namely that of an algebraic weak factorisation system (awfs). We will show that the symmetric effective and the effective Kan fibrations both fit into their own awfs. In the first two subsections, we will recall the definition of a double category and an awfs; the definition of an awfs we give here is the one from [5], where it is called a lifting algebraic weak factorisation system. Indeed, these sections are based on [5] and contain no original work. We will then define a class of double categories based on horn inclusions and study awfss where the left double category comes from this class. We will then characterize the vertical morphisms of the right double category of such an awfs and show that the effective and symmetric effective Kan fibrations satisfy this characterization.